A beam of protons is accelerated through a potential difference of 0.750 kV and then enters a uniform magnetic field traveling perpendicular to the field.

Part A: What magnitude of field is needed to bend these protons in a circular arc of diameter 1.76 m ?

B = ____ T

Part B: What magnetic field would be needed to produce a path with the same diameter if the particles were electrons having the same speed as the protons?

B' = ____ T
Answer #1

PART (A) v = sqrt(2qV/m)

v of proton = sqrt (( 2 x 1.6 x 10^-19 x 750) / (1.67 x 10^-27)) = 379094.414 m/s

r = m v / q B

B = m v / q r

B = (1.67 x 10^-27 x  379094.41) / (1.6 x 10^-19 x .88)

B = 4.49 x 10^-3 T


PART (B)
B' = m v / q r

B' = (9.1 x 10^-31 x 379094.414) / (1.6 x 10^-19 x .88)

B' = 2.45 x 10^-6 T